50g of caco3 is allowed to dissociatedin 22.4 lit vessel at 819 C. if 50% of CaCo3 is left at Equilibrium, active masses of CaCo3, CaO and Co2 are

Grade:7

1 Answers

Chhavi Jain
6 years ago
Dear student, 
 
        vol. of vessel = 22.4 L
       CaCO3 ------>    CaO     +    CO2 
 
t=0    50g or             –                     – 
         50/100= 0.5 moles
 
t=eqbm  0.25 moles       0.25 moles    0.25 moles
 
active mass : CO= 0.25 / 22.4 = 0.011 mol L-1